OpenStudy (anonymous):

A speedboat increases its speed uniformly from 20 m/s to 31 m/s in a distance of 240 m. a) find the magnitude of its acceleration. ___ m/s^2 b) find the time it takes the boat to travel the 240 m distance. ___ s

OpenStudy (anonymous):

A friend sent me this as the answer but I don't quite understand it 240=1/2at^2+V(0)t , V=at+V(0),31=at+20 ,at=11 240=1/2(at)t +20t , 240=5.5t+20t=25.5t, t=240/25.5=9.41sec a=11/9.41=1.16875m/sec^2

OpenStudy (anonymous):

oh dear It is an easy one.. we have v=31m/s u=20m/s d=240m so \[v ^{2}-u ^{2}\]=2 a d; \[31^{2}-20^{2}\]=2xax240; 961-400=2xax240; 561/(2x240)=a so, 561/480=a so,a=1.16m/sec^2 and for the 2nd part, we have,v=u+at putting the values of v u and a we get t; 31=20+1.16t; 31-20=1.16t; 11/1.16=t thus we have t=9.48 sec.

OpenStudy (anonymous):

Thank you, my teachers say I have a problem with over complicating things.

OpenStudy (anonymous):

it was easy isn't it?actually the other guy used the 2nd equation of motion while i used the first and third laws.

OpenStudy (anonymous):

i mean equations